I need help!!
In the following system of linear equations, state how to combine the two equations to obtain one equation with only one variable.
5x - 4y = 6
2x + 3y = 2

Answer by fastblue(13)

5x - 4y = 6
2x + 3y = 2
Take the bottom equation times 5 and the top equation times 2,
and subtract, this will give you -23y =2, which solves to get y=-2/23
Back substitute back in for y, and solve for x, x = 26/23.
Substitute back into original equations to check.